OpenStudy (anonymous):

what about this one Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank).

OpenStudy (nurali):

volume of cylinder=πr^2h Volume of cone=1/3πr^2h=1/3∗15=5 Because r and h are same.

OpenStudy (anonymous):

what about this one Find the volume of a cylinder with a diameter of 8 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank)